a mountain shoots out water that follows a parabolic curve. water comes out of a point on the ground and goes in another port on the ground as well. if the ports are 15 feet above and the water peak at 12 feet from the ground, what is the horizontal distance of the water arc 4 feet from the ground?
Answer by lwsshak3(11628)

a mountain shoots out water that follows a parabolic curve. water comes out of a point on the ground and goes in another port on the ground as well. if the ports are 15 feet above and the water peak at 12 feet from the ground, what is the horizontal distance of the water arc 4 feet from the ground?
***
Place the origin at the peak (vertex) 12 ft above the ground
The equation of the parabola then becomes: y=-Ax^2
solve for A using coordinates (7.5,-12) where the water enters a port on the ground
-12=A(7.5)^2
A=-12/(7.5)^2=-.213
equation: y=-.213x^2
at 4 ft above the ground:
x^2=-4/-.213=18.78
x=4.33 ft
what is the horizontal distance of the water arc 4 feet from the ground? 4.33 ft